# Physics beyond the Standard Model with the DSA-2000

### v1

@@ 2.1 Axion dark matter conversion in NS magnetospheres
It is clear from the expression in Eq. (2) that an ideal target will have a long period $P$ and a strong magnetic field strength $B_{0}$, close to the galactic center (for enhanced DM density), and proximity to Earth. We identify potential NS targets using the pulsar catalog in the Australia Telescope National Facility [71] and the McGill Magnetar Catalog [72] [^3] Online version of the McGill Magnetar Catalog: https://www.physics.mcgill.ca/~pulsar/magnetar/main.html.. Since DSA-2000 is located in Nevada, we consider only NSs with declination above $-30^{\circ}$.
@@
In Fig. 3, we show the expected radio flux density, computed from Eqs. (2.10)-(2), considering targets from the ATNF pulsar catalog [71] and the McGill Magnetar Catalog [72]. We take the representative axion benchmark parameters to be $m_{a}/(2\pi)=1$ GHz and $g_{a\gamma\gamma}=10^{-14}$ GeV-1. We also show the expected flux density from NSs with other values of $P$ and $\dot{P}$ with grey background contours, with $B_{0}$ computed using Eq. (2.5), and fiducial values of $\rho_{\mathrm{DM}}=0.4$ GeV/cm3 and $d=1$ kpc. Note that while increasing $P$ leads to a larger radio flux in general, it also shrinks the conversion radius $r_{c}$ in Eq. (2), which can completely block the axion-photon conversion if $r_{c}$ falls below the NS radius $R_{\mathrm{NS}}$. These NSs, which are colored in black in Fig. 3, will not produce any observable flux.
@@
Additionally, we note that DSA 2000 has the potential of discovering slow magnetars. Indeed, several slow magnetars have been identified in the past few years, including PSR J0901-4046 with $P=75.89$ s [76], GLEAM-X J162759.5-523504.3 with $P=18.2$ min [77], GPM J1839-10 with $P=22.0$ min [78], and ASKAP J1935+2148 with $P=53.8$ min [79]. Among these four magnetars, GPM J1839-10 and ASKAP J1935+2148 are within DSA-2000’s field of view. These slow magnetars also generically have strong surface magnetic fields ($B_{0}\sim 10^{14}-10^{16}$ G), and are thus promising targets for axion search thanks to the $P$ and $B_{0}$ scaling in Eq. (2). We have denoted these slow magnetars as well as the galactic center magnetar with star symbols in the scatter plot in Fig. 3. Due to their long periods and strong magnetic fields, the expected radio flux density from axion-photon conversion can be quite high, ranging from $F\sim 0.01-1$ mJy. We expect DSA-2000 to be able to discover additional slow magnetars with similar physical parameters.
@@
In Fig. 4, we report the projected $5\sigma$ upper limits on $g_{a\gamma\gamma}$ after 10 hours of observation, each on Galactic Center magnetar SGR J1745-2900 [73, 74, 75], and two long-period magnetars, GPM J1839-10 with $P=22.0$ min [78] and ASKAP J1935+2148 with $P=53.8$ min [79] with DSA-2000. The upper limits are derived using Eqs. (2.10)-(2), and setting SNR$=5$ in Eq. (2.14). The parameters for each target are listed in Table 2.
@@
SGR J1745-2900 [73, 74, 75] $3.764$ $2.3\times 10^{14}$ $8.0$ $17^{\text{h}}45^{\text{m}}40^{\text{s}}$ $-29^{\circ}00^{\prime}30^{\prime\prime}$ $0.1$ $1.6\times 10^{5}$
GPM J1839-10 [78] $1318$ $7.0\times 10^{14}$ $5.7$ $18^{\text{h}}39^{\text{m}}2^{\text{s}}$ $-10^{\circ}31^{\prime}49^{\prime\prime}$ $3.7\times 10^{3}$ $2.0$
ASKAP J1935+2148 [79] $3225$ $2.0\times 10^{16}$ $5.0$ $19^{\text{h}}35^{\text{m}}5^{\text{s}}$ $21^{\circ}48^{\prime}42^{\prime\prime}$ $7.0\times 10^{3}$ $0.7$
Table 2: Proposed pulsar and magnetar targets for a DM axion search with DSA-2000. Here, RA and DEC denote the right ascension and declination of the star, respectively. The values of $R_{\mathrm{GC}}$ for GPM J1839-10 and ASKAP J1935+2148 are inferred using $d$, RA, and DEC, while the value of $R_{\mathrm{GC}}$ for SGR J1745-2900 is taken from Ref. [75]. The value of $\rho_{\mathrm{DM}}$ for each star is computed assuming an NFW halo profile (cf. Eq. (2.9)).
@@
Figure 4: Projected $5\sigma$ upper limits on $g_{a\gamma\gamma}$ are shown assuming 10 hours of observation time ($\Delta t_{\mathrm{obs}}=10$ hrs) on the Galactic Center magnetar SGR J1745-2900 [73, 74, 75] (red dashed), and two long-period magnetars: GPM J1839-10 with $P=22.0$ min [78] (orange solid) and ASKAP J1935+2148 with $P=53.8$ min [79] (green solid), observed with DSA 2000. The properties of each target are listed in Table 2. Here, we assume $M_{\mathrm{NS}}=1M_{\odot}$, $R_{\mathrm{NS}}=10$ km, $\theta=120^{\circ}$, $\theta_{m}=10^{\circ}$, and $\mathrm{SEFD}=2.5$ Jy. The DM density near the NS is estimated using an NFW profile with a scale radius of $r_{s}=8$ kpc, with the Galactic Center magnetar’s distance from the Galactic Center taken to be $R_{\mathrm{GC}}=0.1$ pc. Existing constraints from ADMX [80, 81, 82, 83, 84, 85, 86], RBF [87, 88], UF [89, 90], CAPP [91, 92, 93, 94, 95, 96, 97, 98, 99, 100], CAST [101, 102, 103], globular clusters [104, 105], pulsar polar cap [106], solar neutrinos [107], MWD [108, 109], and MeerKAT [19] are shaded in grey (see Ref. [110]). The QCD axion parameter space is shaded in yellow [111].
@@
We find that DSA-2000 will be sensitive to axions with masses in the range $2.9\,\mathrm{\mu eV}\lesssim m_{a}\lesssim 8.3\,\mathrm{\mu eV}$ and coupling strengths $g_{a\gamma\gamma}\gtrsim 10^{-13}\,\mathrm{GeV}^{-1}$, assuming axions constitute all of DM. Part of the axion mass range within DSA-2000’s sensitivity has already been ruled out by a combination of direct detection and astrophysical constraints. However, DSA-2000 can place constraints on some previously unexplored regions of parameter space, such as at $m_{a}\approx 4.2\,\mathrm{\mu eV},5.6\,\mathrm{\mu eV},6.4\,\mathrm{\mu eV}$ and $7.2\,\mathrm{\mu eV}$. It is also interesting to note that both SGR J1745-2900 and ASKAP J1935+2148 are projected to place almost identical upper limits on $g_{a\gamma\gamma}$ in a log-log plot. This is because the photon flux density measured on Earth scales parametrically as $\propto B_{0}^{5/6}P^{7/6}\rho_{\mathrm{DM}}^{\infty}d^{-2}$ (cf. Eqs. (2)-(2)), for which these two stars yield numerical values within $\sim 20\%$ of each other based on their parameters listed in Table 2.
@@
Our analysis is based on a number of simplifying assumptions. We model axion-photon mixing using a 1D approximation in which photon trajectories are radial such as in [6]. More detailed computations that account for, e.g., 3D effects, non-radial axion and photon trajectories, mode mixing, small scale plasma fluctuations, have been performed in [56, 57, 68, 20, 112]. The numerical analysis of axion-photon mixing performed in [112] agrees well with the results of [68], on which our calculations are based. An additional complication in modeling magnetar magnetospheres is the presence of magnetic “twist” – that is, regions where ×$\boldsymbol{\nabla}\times\mathbf{B}\neq 0$. Such twist generates strong currents in the magnetosphere, supported by plasma with densities far exceeding the GJ value. The implications of twisted magnetospheres were first explored in [20], with a more detailed analysis presented in [27]. The latter study found that in certain regimes—particularly as the magnetar enters quiescence—large regions of the magnetosphere may resemble the untwisted, GJ-like configuration. This behavior supports the estimates used in our modeling, but more observations of the considered sources are required to reduce systematic uncertainties.
@@ 2.2 The need for non-dark matter axion probes
A major obstacle in the way of axion DM detection is the theoretical uncertainty in predicting the present-day axion abundance. For instance, the cosmology of the QCD axion depends critically on when PQ symmetry is broken relative to inflationary reheating. For pre-inflationary symmetry breaking, the present-day axion abundance is degenerate with the initial misalignment angle, providing little information on the axion mass [37]. For post-inflationary symmetry breaking, there is a unique axion mass that provides the observed DM abundance, but determining that mass has been the subject of decades of intensive numerical simulations [113, 38, 114, 115, 39, 116, 41, 117, 118, 119, 111, 120, 121]. Despite disagreements on the determination of the axion mass by different groups, most predict the axion mass to be $m_{a}\gtrsim\mu$eV. Moreover, it is possible that the axion mass lives outside the aforementioned range and makes up only a subcomponent of DM, reducing the sensitivity of probes that rely on axions making up all of DM.
Apart from being a compelling DM candidate, the axion holds significant theoretical interest in its own right. As discussed in the beginning of Sec. 2, the QCD axion provides an elegant solution to the strong CP problem in the Standard Model, releasing the fine-tuning constraint on the CP-violating parameter $\bar{\Theta}$. Additionally, in string theory, axions arise naturally in the compactifications of Calabi-Yau manifolds, and the detection of a QCD axion could facilitate meaningful statements about the string theory landscape [122]. Therefore, the discovery of an axion, especially the QCD axion, could have serious theoretical implications even if it does not constitute a significant fraction of DM. Such opportunities motivate Light-Shining-through-Wall (LSW) experiments where axions are generated from coupling to SM fields, such as $\mathcal{L}_{a}\supset-g_{a\gamma\gamma}a{\bf E}\cdot{\bf B}$. Unlike in laboratories where the magnetic field is well controlled but weak, astrophysical systems like NSs offer unique advantages with their strong magnetic fields. In the following subsection, we explore signals from NS-originated axion clouds converting to photons through the NS magnetosphere.
@@ 2.3 Vacuum gaps around neutron stars
The GJ model describes the minimum charge density required to screen the rotation-induced electric field component, $E_{\parallel}$, parallel to the background magnetic field. Crucially, this model is insufficient to describe regions that do not co-rotate with the neutron star. Many important deviations from the GJ model are reviewed below. The GJ model breaks down near the light cylinder (LC), a cylinder that is coaxial with the NS rotation axis and with radius $R_{\rm LC}=c/\Omega$, where $\Omega$ is the NS rotation frequency. The LC defines the boundary beyond which co-rotation requires superluminal propagation. In the equatorial region near the LC, the GJ density diverges, signaling a breakdown of the model. The LC divides the region surrounding the NS into the “closed” and “open” magnetospheres illustrated in Fig. 5, where the former contains all magnetic field lines whose maximum extent from the NS lies within the magnetosphere (thus allowing the field lines to close) and the latter contains field lines that exit the LC and whose closure within the magnetosphere is forbidden by causality. Application of the GJ model is appropriate only in the closed magnetosphere and well within the LC. The footpoints of all open field lines form a region on the NS surface called the polar cap (PC), represented as cyan cylinders in Fig. 5. For a dipolar magnetic field, the PC is described by the region on the NS surface with $\theta\leq\theta_{\rm pc}\equiv\sqrt{\Omega R_{\rm NS}}$, where $\theta$ is the polar angle in spherical coordinates (with $\theta=0$ corresponding to the magnetic pole, which here and below is assumed to be aligned with the rotation axis, such that $\theta_{m}=0$). [^4] This expression may be derived by noticing that dipolar field lines are contours of fixed $r/\sin^{2}(\theta)$. Each contour may be defined by the polar angle of its footpoint, $\theta_{f}$, such that $r/\sin^{2}(\theta)=R_{\rm NS}/\sin^{2}(\theta_{f})$. The maximum radius a field line extends from the NS occurs when $\theta=\pi/2$, giving $R_{\rm max}=R_{\rm NS}/\sin^{2}(\theta_{f})$. The last closed field line, which demarcates the edge of the polar cap has maximum radius, $R_{\rm max}=R_{\rm LC}$, which gives the quoted expression for $\theta_{\rm pc}$ in the limit that $\Omega R_{\rm NS}\ll 1$, which is always the case for pulsars of interest. Particles flowing along open field lines are evacuated from the polar region, leading to the formation of a vacuum gap right above the PC. The dearth of pairs in the gap leads to a large longitudinal voltage drop along the gap, which grows until it becomes unstable to runaway $e^{\pm}$ pair creation, screening the gap [123]. The competition between pair evacuation along open field lines and screening leads to a rich spatio-temporal evolution of $E_{\parallel}$ (see right panel in Fig. 5), with important implications for axion production as we describe in the following subsection  [124, 106].

### v2

@@ 2.1 Axion dark matter conversion in neutron star magnetospheres
It is clear from the expression in Eq. (2) that an ideal target will have a long period $P$ and a strong magnetic field strength $B_{0}$, close to the galactic center (for enhanced DM density), and proximity to Earth. We identify potential NS targets using the pulsar catalog in the Australia Telescope National Facility [Manchester:2004bp] and the McGill Magnetar Catalog [Olausen:2013bpa] [^2] Online version of the McGill Magnetar Catalog: https://www.physics.mcgill.ca/~pulsar/magnetar/main.html.. Since DSA-2000 is located in Nevada, we consider only NSs with declination above $30^{\circ}$.
@@
In Fig. 3, we show the expected radio flux density, computed from Eqs. (2.10)-(2), considering targets from the ATNF pulsar catalog [Manchester:2004bp] and the McGill Magnetar Catalog [Olausen:2013bpa]. We take the representative axion benchmark parameters to be $m_{a}/(2\pi)=1$ GHz and $g_{a\gamma\gamma}=10^{-14}$ GeV-1. We also show the expected flux density from NSs with other values of $P$ and $\dot{P}$ with grey background contours, with $B_{0}$ computed using Eq. (2.5), and fiducial values of $\rho_{\mathrm{DM}}=0.4$ GeV/cm3 and $d=1$ kpc. Note that while increasing $P$ leads to a larger radio flux in general, it also shrinks the conversion radius $r_{c}$ in Eq. (2), which can completely block the axion-photon conversion if $r_{c}$ falls below the NS radius $R_{\mathrm{NS}}$. These NSs, which are colored in black in Fig. 3, will not produce any observable flux.
@@
Additionally, we note that DSA 2000 has the potential of discovering slow magnetars. Indeed, several slow magnetars have been identified in the past few years, including PSR J0901-4046 with $P=75.89$ s [Caleb:2022xyo], GLEAM-X J162759.5-523504.3 with $P=18.2$ min [Hurley-Walker:2022], GPM J1839-10 with $P=22.0$ min [Hurley-Walker:2023], and ASKAP J1935+2148 with $P=53.8$ min [Caleb:2024nrq]. Among these four magnetars, GPM J1839-10 and ASKAP J1935+2148 are within DSA-2000’s field of view. These slow magnetars also generically have strong surface magnetic fields ($B_{0}\sim 10^{14}-10^{16}$ G), and are thus promising targets for axion search thanks to the $P$ and $B_{0}$ scaling in Eq. (2). We have denoted these slow magnetars as well as the galactic center magnetar with star symbols in the scatter plot in Fig. 3. Due to their long periods and strong magnetic fields, the expected radio flux density from axion-photon conversion can be quite high, ranging from $F\sim 0.01-1$ mJy. We expect DSA-2000 to be able to discover additional slow magnetars with similar physical parameters.
@@
In Fig. 4, we report the projected $5\sigma$ upper limits on $g_{a\gamma\gamma}$ after 10 hours of observation, each on Galactic Center magnetar SGR J1745-2900 [Kennea:2013dfa, Mori:2013yda, Shannon:2013hla], and two long-period magnetars, GPM J1839-10 with $P=22.0$ min [Hurley-Walker:2023] and ASKAP J1935+2148 with $P=53.8$ min [Caleb:2024nrq] with DSA-2000. For comparison, we also show upper limits from an ordinary (i.e. non-magnetar) pulsar J0250+5854, with $P=23.5$ s [Tan:2018rhg], which is the second-slowest-spinning radio pulsar known to date. [^3] The slowest-spinning radio pulsar known to date is PSR J0901-4046 with $P=75.9$ s [Caleb:2022xyo], which is one of the long-period magnetars mentioned in the previous paragraph. It is not within DSA-2000’s field of view. The upper limits are derived using Eqs. (2.10)-(2), and setting SNR=5=5 in Eq. (2.14). The parameters for each target are listed in Table 2.
@@
SGR J1745-2900 [Kennea:2013dfa, Mori:2013yda, Shannon:2013hla] $3.764$ $2.3\times 10^{14}$ $8.0$ $17^{\text{h}}45^{\text{m}}40^{\text{s}}$ $29^{\circ}00^{\prime}30^{\prime\prime}$ $0.1$ $1.6\times 10^{5}$
GPM J1839-10 [Hurley-Walker:2023] $1318$ $7.0\times 10^{14}$ $5.7$ $18^{\text{h}}39^{\text{m}}2^{\text{s}}$ $10^{\circ}31^{\prime}49^{\prime\prime}$ $3.7\times 10^{3}$ $2.0$
ASKAP J1935+2148 [Caleb:2024nrq] $3225$ $2.0\times 10^{16}$ $5.0$ $19^{\text{h}}35^{\text{m}}5^{\text{s}}$ $21^{\circ}48^{\prime}42^{\prime\prime}$ $\times 10^{3} 0.70.7$
J0250+5854 [Tan:2018rhg] $23.54$ $2.6\times 10^{13}$ $1.6$ $02^{\text{h}}50^{\text{m}}18^{\text{s}}$ $58^{\circ}54^{\prime}01^{\prime\prime}$ $9.7\times 10^{3}$ $0.3$
Table 2: Proposed pulsar and magnetar targets for a DM axion search with DSA-2000. Here, RA and DEC denote the right ascension and declination of the star, respectively. The values of $R_{\mathrm{GC}}$ for GPM J1839-10 and ASKAP J1935+2148 are inferred using $d$, RA, and DEC, while the value of $R_{\mathrm{GC}}$ for SGR J1745-2900 is taken from Ref. [Shannon:2013hla]. The value of $\rho_{\mathrm{DM}}$ for each star is computed assuming an NFW halo profile (cf. Eq. (2.9)).
@@
Figure 4: Projected $5\sigma$ upper limits on $g_{a\gamma\gamma}$ are shown assuming 10 hours of observation time ($\Delta t_{\mathrm{obs}}=10$ hrs) on the Galactic Center magnetar SGR J1745-2900 [Kennea:2013dfa, Mori:2013yda, Shannon:2013hla] (red dashed), two long-period magnetars: GPM J1839-10 with $P=22.0$ min [Hurley-Walker:2023] (orange solid) and ASKAP J1935+2148 with $P=53.8$ min [Caleb:2024nrq] (green solid), and an ordinary pulsar J0250+5854 with $P=23.5$ s [Tan:2018rhg] (blue solid), observed with DSA 2000. The properties of each target are listed in Table 2. Here, we assume $M_{\mathrm{NS}}=1M_{\odot}$, $R_{\mathrm{NS}}=10$ km, $\theta=120^{\circ}$, $\theta_{m}=10^{\circ}$, and $\mathrm{SEFD}=2.5$ Jy. The DM density near the NS is estimated using an NFW profile with a scale radius of $r_{s}=8$ kpc, with the Galactic Center magnetar’s distance from the Galactic Center taken to be $R_{\mathrm{GC}}=0.1$ pc. Existing constraints from ADMX [Asztalos_2010, ADMX:2018gho, ADMX:2018ogs, ADMX:2019uok, ADMX:2021mio, ADMX:2021nhd, ADMX:2024xbv], RBF [PhysRevLett.59.839, Wuensch:1989sa], UF [PhysRevD.42.1297, Hagmann:1996qd], CAPP [Lee:2020cfj, Jeong:2020cwz, CAPP:2020utb, Lee:2022mnc, Yoon:2022gzp, Kim:2022hmg, Yi:2022fmn, Yang:2023yry, Kim:2023vpo, CAPP:2024dtx], CAST [CAST:2007jps, CAST:2017uph, CAST:2024eil], globular clusters [Ayala:2014pea, Dolan:2022kul], pulsar polar cap [noordhuis2023novel], solar neutrinos [Vinyoles_2015], MWD [Dessert:2021bkv, Dessert:2022yqq, Benabou:2025jcv], and MeerKAT [Battye:2023oac] are shaded in grey (see Ref. [AxionLimits]). The QCD axion parameter space is shaded in yellow [Saikawa:2024bta].
@@
We find that DSA-2000 will be sensitive to axions with masses in the range $2.9\,\mathrm{\mu eV}\lesssim m_{a}\lesssim 8.3\,\mathrm{\mu eV}$ and coupling strengths $g_{a\gamma\gamma}\gtrsim 2\times 10^{-13}\,\mathrm{GeV}^{-1}$, assuming axions constitute all of DM. However, the axion mass range within DSA-2000’s sensitivity has already been ruled out by a combination of direct detection and astrophysical constraints. As expected, magnetars can generally place tighter constraints on the axion coupling strength thanks to their strong magnetic field, although ordinary pulsars are subjected to less modeling uncertainties. It is also interesting to note that both SGR J1745-2900 and ASKAP J1935+2148 are projected to place almost identical upper limits on $g_{a\gamma\gamma}$ in a log-log plot. This is because the photon flux density measured on Earth scales parametrically as $\propto B_{0}^{5/6}P^{7/6}\rho_{\mathrm{DM}}^{\infty}d^{-2}$ (cf. Eqs. (2)-(2)), for which these two stars yield numerical values within $\sim 20\%$ of each other based on their parameters listed in Table 2.
@@
Our analysis is based on a number of simplifying assumptions. We model axion-photon mixing using a 1D approximation in which photon trajectories are radial such as in [Hook:2018iia]. More detailed computations that account for, e.g., 3D effects, non-radial axion and photon trajectories, mode mixing, small scale plasma fluctuations, have been performed in [Witte2021, millar2021axionphotonUPDATED, McDonald:2023ohd, McDonald:2023shx, Gines:2024ekm]. The numerical analysis of axion-photon mixing performed in [Gines:2024ekm] agrees well with the results of [McDonald:2023ohd], on which our calculations are based. We also neglect Euler–Heisenberg corrections to the photon dispersion relation, which are expected to become relevant for relativistic axions [Long:2024qvd]. This omission is justified in our case, as the axions we consider are at most semi-relativistic. An additional complication in modeling magnetar magnetospheres is the presence of magnetic “twist” – that is, regions where $\boldsymbol{\nabla}\times\mathbf{B}\neq 0$. Such twist generates strong currents in the magnetosphere, supported by plasma with densities far exceeding the GJ value. The implications of twisted magnetospheres were first explored in [McDonald:2023shx], with a more detailed analysis presented in [Roy:2025mqw]. The latter study found that in certain regimes—particularly as the magnetar enters quiescence—large regions of the magnetosphere may resemble the untwisted, GJ-like configuration. This behavior supports the estimates used in our modeling, but more observations of the considered sources are required to reduce systematic uncertainties.
@@ 2.2 The need for non-dark matter axion probes
[…]
Apart from being a compelling DM candidate, the axion holds significant theoretical interest in its own right. As discussed in the beginning of Sec. 2, the QCD axion provides an elegant solution to the strong CP problem in the Standard Model, releasing the fine-tuning constraint on the CP-violating parameter $\bar{\Theta}$. Additionally, in string theory, axions arise naturally in the compactifications of Calabi-Yau manifolds, and the detection of a QCD axion could facilitate meaningful statements about the string theory landscape [Gendler:2024adn]. Therefore, the discovery of an axion, especially the QCD axion, could have serious theoretical implications even if it does not constitute a significant fraction of DM. Such opportunities motivate Light-Shining-through-Wall (LSW) experiments where axions are generated from coupling to SM fields, such as $\mathcal{L}_{a}\supset-g_{a\gamma\gamma}a{\bf E}\cdot{\bf B}$. Unlike in laboratories where the magnetic field is well controlled but weak, astrophysical systems like NSs offer unique advantages with their strong magnetic fields. In the following subsection, we explore signals from NS-originated axion clouds converting to photons through the NS magnetosphere.
@@ 2.3 Vacuum gaps around neutron stars
The GJ model describes the minimum charge density required to screen the rotation-induced electric field component, $E_{\parallel}$, parallel to the background magnetic field. Crucially, this model is insufficient to describe regions that do not co-rotate with the neutron star. Many important deviations from the GJ model are reviewed below. The GJ model breaks down near the light cylinder (LC), a cylinder that is coaxial with the NS rotation axis and with radius $R_{\rm LC}=c/\Omega$, where $\Omega$ is the NS rotation frequency. The LC defines the boundary beyond which co-rotation requires superluminal propagation. In the equatorial region near the LC, the GJ density diverges, signaling a breakdown of the model. The LC divides the region surrounding the NS into the “closed” and “open” magnetospheres illustrated in Fig. 5, where the former contains all magnetic field lines whose maximum extent from the NS lies within the magnetosphere (thus allowing the field lines to close) and the latter contains field lines that exit the LC and whose closure within the magnetosphere is forbidden by causality. Application of the GJ model is appropriate only in the closed magnetosphere and well within the LC. The footpoints of all open field lines form a region on the NS surface called the polar cap (PC), represented as cyan cylinders in Fig. 5. For a dipolar magnetic field, the PC is described by the region on the NS surface with $\theta\leq\theta_{\rm pc}\equiv\sqrt{\Omega R_{\rm NS}}$, where $\theta$ is the polar angle in spherical coordinates (with $\theta=0$ corresponding to the magnetic pole, which here and below is assumed to be aligned with the rotation axis, such that $\theta_{m}=0$). [^4] This expression may be derived by noticing that dipolar field lines are contours of fixed $r/\sin^{2}(\theta)$. Each contour may be defined by the polar angle of its footpoint, $\theta_{f}$, such that $r/\sin^{2}(\theta)=R_{\rm NS}/\sin^{2}(\theta_{f})$. The maximum radius a field line extends from the NS occurs when $\theta=\pi/2$, giving $R_{\rm max}=R_{\rm NS}/\sin^{2}(\theta_{f})$. The last closed field line, which demarcates the edge of the polar cap has maximum radius, $R_{\rm max}=R_{\rm LC}$, which gives the quoted expression for $\theta_{\rm pc}$ in the limit that $\Omega R_{\rm NS}\ll 1$, which is always the case for pulsars of interest. Particles flowing along open field lines are evacuated from the polar region, leading to the formation of a vacuum gap right above the PC. The dearth of pairs in the gap leads to a large longitudinal voltage drop along the gap, which grows until it becomes unstable to runaway $e^{\pm}$ pair creation, screening the gap [1971ApJ...164..529S]. The competition between pair evacuation along open field lines and screening leads to a rich spatio-temporal evolution of $E_{\parallel}$ (see right panel in Fig. 5), with important implications for axion production as we describe in the following subsection  [Prabhu:2021zve, Noordhuis:2022ljw].